OpenStudy (anonymous):

y=3x^3-3x find the zero and the state the multiplicity of multiple zeros

OpenStudy (anonymous):

What exactly do you want us to do?

OpenStudy (anonymous):

@Riddellikins find the zero and the state the multiplicity of multiple zeros

OpenStudy (anonymous):

0 = 3x^3 - 3x 0 = 3x^2 - 3 (divide both sides by x) 3x^2 = 3 x^2 = 1 x = 1, -1 1 and -1 are the two zeros I don't know what "the multiplicity of multiple zeros" means